The volume, L liters, of paint in a plastic tub may be assumed to be normally distributed with mean 10.25 and variance σ^2.
(a) assuming that variance = 0.04, determine P(L<10).
(b) Find the value of standard deviation so that 98% of tubs contain more than 10 liters of paint.
1
Expert's answer
2019-09-11T13:28:58-0400

Suppose "L" is a variable representing the volume of paint in a plastic tub.

"L" is normally distributed with mean "\mu=10.25" and variance "\sigma^2:L\sim N(10.25, \sigma^2)."

If "L\sim(N, \sigma^2)," then "Z={L-\mu \over \sigma}\sim N(0, 1)"

(a) Assume that "\sigma^2=0.04"


"Z={L-\mu \over \sigma}={10-10.25 \over \sqrt{0.04}}=-1.25"

Then


"P(L<10)=P(Z<-1.25)=0.105650""P(L<10)=0.105650"


(b) Find the value of standard deviation so that 98% of tubs contain more than 10 liters of paint.

Given that


"P(L>10)=0.98"

Then


"P(L>10)=P(Z>Z^*)=0.98=>""=>Z^*=-2.053749"

"{L-\mu \over \sigma}=Z^*""\sigma={L-\mu \over Z^*}"

"\sigma={10-10.25 \over -2.053749}\approx0.1217\approx \sqrt{0.0148}"

If the value of standard deviation is "0.1217," then 98% of tubs contain more than 10 liters of paint.